Claim: Fewer than 98% of adults have a cell phone. In a reputable poll of 1200 adults, 90% said that they have a cell phone. Find the value of the test statistic. 


The value of the test statistic is _______ 

(Round to two decimal places as needed.)
Answer #1

Solution :

The null and alternative hypothesis is

H0 : p = 0.98

Ha : p < 0.98

р = 0.90

1 - P0 = 1 - 0.98 = 0.02

Test statistic = z

= р - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.90 - 0.98 / [\sqrt(0.98 * 0.02) / 1200]

= - 0.08 / 0.989941

= - 0.08

The value of test statistic is - 0.08